4. a. Assume Be,-50 percent and ps-2 percent. Calculate ścPs/s) and šs/ps) according to the precise and approximate dynamic PPP conditions. US Does empirical evidence suggest PPP may be more likely to hold in the short run or the long run? Can you suggest an explanation for why this might be true?
Answer #1

Question 4 (a)

According to the precise dynamic PPP conditions,

S (PS / $) = (PMex - PUS) / (1 + PUS) = (50% - 2%) / (1 + 2%) = 0.4706 = 47.06%

S ($ / PS) = (PUS - PMex) / (1 + PMex) = (2% - 50%) / (1 + 50%) = -0.3200 = -32.00%

According to approximate dynamic PPP conditions,

S (PS / $) = (PMex - PUS) = (50% - 2%) = 48% = 0.48

S ($ / PS) = (PUS - PMex) = (2% - 50%) = -48% = -0.48

----------------

Next Question

Deviations have been observed from PPP in the short run as well as long run. The question of long run is still unanswered, that it, when does the long run arrive.

However, the deviations from PPP had been relatively lower in the long run than in the short run. In the short run, there are always deviations from the PPP.

Empirical studies and evidences support the argument above and hence they suggest PPP may be more likely to hold ture in the long run.

In order to provide explanation, we will have to go behind the assumptions leading to PPP and check whether those assumptions are valid more in long run or short run:

  • One assumption in PPP is that prices of goods and services are predominant driver of exchange rate. In reality, there are many other factors like interest rate, costs, political stability, economic stability, etc that also impact the exchange rate. In the short run these factors may dominate over the effect of inflation captured in the cost of goods and services and hence exchange rate may deviate from PPP. In the long run, these other factors usually don't dominate the impact of inflation on exchange rate. hence deviation from PPP in the long run is lower.
  • Another assumption in PPP is that law of one price holds true. Studies have shown that law of one price appears to hold true in long run but not in short run.